# A Scalable Optimization Approach for Equitable Facility Location: Methodology and Transportation Applications

### v3

@@ 5 Discussion and Conclusions
This research addresses a fundamental challenge in transportation facility location: how to optimize the Kolm-Pollak EDE while maintaining computational tractability for real-world networks. We have demonstrated that our approach not only scales to very large problems but achieves this without sacrificing solution quality. The linear proxy formulation (KPL) performs comparably to the p-median model in computational time while delivering solutions that better serve both average and worst-case users.
@@
    The approach is computationally efficient: The (KPL) model solved instances with over 200 million binary variables, demonstrating comparable performance to the p-median model and significantly outperforming the p-center approach. This represents a significant improvement over previous equity-based approaches that became intractable for realistic problem sizes.
@@
    Equality can be improved without compromising the average travel time: As shown in Section 4.4.1, the model balances efficiency and equity. On average across all cities, sacrificing less than 10 meters in mean travel distance yielded improvements of over 400 meters in maximum distance. This demonstrates that incorporating equity considerations need not compromise overall system efficiency.

### v4

@@ 6 Discussion and Conclusions
This research addresses a fundamental challenge in transportation facility location: how to optimize the Kolm-Pollak EDE while maintaining computational tractability for real-world networks. We have demonstrated that our approach not only scales to very large problems but achieves this without sacrificing solution quality. The linear proxy formulation (KPL) performs comparably to the p-median model computationally while delivering solutions that better serve both average and worst-case users. Similar to DOMP models, which do not yet scale computationally, (KPL) delivers nuanced optimal solutions, equitably distributing travel burdens within the interior of a distribution of distances, while simultaneously achieving near-optimal average welfare.
@@
    The approach is computationally efficient: The (KPL) model solved instances with over 200 million binary variables, demonstrating comparable performance to the p-median model and significantly outperforming the p-center and p-centdian approaches. This represents a significant improvement over previous equity-based approaches that became intractable for realistic problem sizes.
@@
    Equality can be improved without compromising the average travel time: As shown in Section 5.3, the model balances efficiency and equity. On average across all cities, sacrificing less than 10 meters in mean travel distance yielded improvements of over 400 meters in maximum distance. This demonstrates that incorporating equity considerations need not compromise overall system efficiency.
@@ 7 Acknowledgements
Appendix A Theorem and Proposition Proofs
Proposition 1.
Proof.
$=p_{r}e^{-\kappa d_{r,s^{\prime}}}$
$=\sum_{s\in S}p_{r}y_{r,s}e^{-\kappa d_{r,s}}.$
Theorem 3.
$\rho~:=~Te^{-\kappa\mathcal{K}}(e^{-\kappa\sigma}-1),$
Proof.
Theorem 4.
$\hat{E}(\Delta,\sigma^{*})~:=~\hat{\sigma}-\sigma^{*}~=~-\frac{1}{\kappa}\ln\left(e^{\kappa\sigma^{*}}+e^{-\kappa\Delta}(1-e^{\kappa\sigma^{*}})\right).$
Proof.
[…]
so that,
Theorem 5.
Suppose $\sigma^{*}>0$ is fixed:
  1. (1)
  2. (2)
  3. (3)
Suppose $\Delta\in\mathbb{R}$ is fixed:
  1. 4.
Proof.
(1) Suppose $\sigma^{*}>0$ is fixed. We have
Recalling that $\kappa<0$, we have that $e^{-\kappa\sigma^{*}}>1$, so that $\frac{\delta}{\delta\Delta}\hat{E}(\Delta,\sigma^{*})\in(0,1)$. Thus, $\hat{E}(\Delta,\sigma^{*})$ is increasing (though not quickly) in $\Delta$. It is straightforward to evaluate $\hat{E}(0,\sigma^{*})=0$. Thus, when $\Delta>0$, $\hat{E}(\Delta,\sigma^{*})>0$ and $\hat{E}(\Delta,\sigma^{*})=|\hat{E}(\Delta,\sigma^{*})|$ is increasing in $|\Delta|$. Similarly, when $\Delta<0$, $\hat{E}(0,\sigma^{*})<0$ and $\hat{E}(\Delta,\sigma^{*})=|\hat{E}(\Delta,\sigma^{*})|$ is increasing in $|\Delta|$.
Proposition 8.
$\mathcal{K}^{all}~\leq~\mathcal{K}^{*}~\leq~\mathcal{K}^{rem}.$
Proof.
Theorem 10.
Proof.
Thus,
Finally,
$~=~\mathcal{K}^{*}+\sigma^{*}+\Delta$
$~\leq~\mathcal{K}^{*}+\sigma^{*}+\hat{E}(\Delta,\sigma^{*})$
$~\leq~\mathcal{K}^{all}+\sigma^{all},$
which means that $\sigma^{*}\leq\sigma^{all}$, as required. ∎
Proposition 11.
Proof.
Due to constraints (1e), for all feasible $\mathbf{x}:=(x_{1},x_{2},\dots,x_{|S|})$, $q=-\kappa\sum_{s\in U}cx_{s}$ can only take the values of the selected linearization points. At these values of $q$, the approximating tangent lines match the exponential expression, $e^{q}$, exactly. ∎
Proposition 12.
Proof.
Proposition 14.
Proof.
where the final inequality is due to the Cauchy-Schwarz inequality. ∎
Theorem 16.
where $A(w):=e^{\frac{we^{w}}{e^{w}-1}-1}-\frac{we^{w}}{e^{w}-1}$.
Proof.
$~\geq~0.$
We have $\ddot{\sigma}\leq\hat{\sigma}$ because $err(q^{*},w)\geq 0$, and for Equation (18),
For the first part of (19), if $\Delta\leq 0$, then $1+e^{\kappa\sigma^{*}}(e^{\kappa\Delta}-1)\geq 1$, so that $\hat{\sigma}-\ddot{\sigma}\leq\frac{1}{\kappa}\ln\left(1-A(w)\right)$. For the second part of (19), if $\hat{\mathcal{K}}=\mathcal{K}^{rem}$, then $\mathcal{K}^{*}+\sigma^{*}\leq\mathcal{K}^{*}+\hat{\sigma}\leq\mathcal{K}^{rem}+0$, where the first inequality is due to Theorem 5 and the second is due to the optimality of (KPLp). Thus, $0\leq\sigma^{*}\leq\mathcal{K}^{rem}-\mathcal{K}^{*}=\Delta$, and we have,
$1+e^{\kappa\sigma^{*}}(e^{\kappa\Delta}-1)~\geq~1+e^{\kappa\Delta}(e^{\kappa\Delta}-1)~\geq~0.75.$